# LLMs as Planning Modelers: A Survey for Leveraging Large Language Models to Construct Automated Planning Models

### v1

@@ Abstract
Large Language Models (LLMs) excel in various natural language tasks but often struggle with long-horizon planning problems requiring structured reasoning. This limitation has drawn interest in integrating neuro-symbolic approaches within the Automated Planning (AP) and Natural Language Processing (NLP) communities. However, identifying optimal AP deployment frameworks can be daunting. This paper aims to provide a timely survey of the current research with an in-depth analysis, positioning LLMs as tools for extracting and refining planning models to support reliable AP planners. By systematically reviewing the current state of research, we highlight methodologies, and identify critical challenges and future directions, hoping to contribute to the joint research on NLP and Automated Planning.
@@ 1 Introduction
Large Language Models (LLMs) has marked a significant paradigm shift in AI, sparking claims regarding emergent reasoning capabilities within LLMs (Wei et al. 2022a) and their potential integration into automated planning for agents (Pallagani et al. 2023). While LLMs, due to the prowess of distributed representation and learning, excel at System I tasks, planning—an essential aspect of System II cognition (Daniel 2017)—remains to be a significant bottleneck (Bengio 2020). Furthermore, LLMs face challenges with long-term planning and reasoning, often producing unreliable plans (Valmeekam, Stechly, and Kambhampati 2024; Pallagani et al. 2023; Momennejad et al. 2023), frequently failing to account for the effects and requirements of actions as they scale (Stechly, Valmeekam, and Kambhampati 2024), and their performance degrades with self-iterative feedback (Stechly, Marquez, and Kambhampati 2023; Valmeekam, Marquez, and Kambhampati 2023; Huang et al. 2024a).
@@
Figure 1: Distinction of LLM-planning: (a) uses LLMs for direct planning; (b) uses LLMs to generate AP specifications for existing task planning methods such as PDDL.
Acknowledging the challenges of LLMs’ direct planning capabilities, Automated Planning (AP), or AI planning, presents a promising alternative to generate robust, optimal plans through logical and computational methods. Meanwhile, LLMs excel at extracting and refining classical planning models from Natural Language (NL), leaving the plan generation to classical planners and heuristics.
@@
This paper is driven by the fragmented landscape in the current literature, where many surveys lack a cohesive overview of LLM integration within this field. Our focus stems from the need to address these gaps and provide a clear framework, highlighting the importance of aligning LLM capabilities with areas where they offer tangible benefits. The motivation for this approach is threefold: (i) Planning Accuracy: LLMs can help ensure that all relevant factors are considered, reducing the risk of overlooked constraints (Huang, Lipovetzky, and Cohn 2024). (ii) Adaptability: LLMs can aid systems in adapting to dynamic environments to capture real-world nuances better, reducing the need for manual edits (Dagan, Keller, and Lascarides 2023). (iii) Agnostic Modeling: LLMs enable domain-agnostic models reducing reliance on specialized expertise, a major bottleneck in AP (Gestrin, Kuhlmann, and Seipp 2024).
@@
Our survey’s taxonomy is divided into three key areas: Model Generation, further consisting of (i) Task Generation (sec. 3.1.1), which translates initial and goal states along with object assignments; (ii) Domain Generation (sec. 3.1.2), constructing predicates and action schemas; and (iii) Hybrid Generation (sec. 3.1.3), encapsulating both task instance and domain generation. Model refinement and repair are addressed in Model Editing (sec. 3.2), while LLM-AP benchmarks are discussed in Model Benchmarks (sec. 3.3).
@@
To our knowledge, this is the first comprehensive survey of LLM-driven AP-model construction. Our contributions can be summarized as follows:
@@
  • A critical survey of LLM-driven Automated Planning (AP) model generation, editing, and AP-LLM benchmarks, structured within our taxonomy.
@@ 2.3 Large Language Models + Planning
Large Language Models (LLMs) have shown promise in generating highly structured outputs from natural language (NL) descriptions. Frameworks like StarCoder (Li et al. 2023), CodeGen (Nijkamp et al. 2023) and PLANSEARCH (Wang et al. 2024) have demonstrated LLMs’ ability to translate NL to executable code. Previous work has translated formal intermediate representations such as Temporal Logic (Chen et al. 2023b; Pan, Chou, and Berenson 2023; Mao et al. 2024; Cosler et al. 2023; Liu et al. 2023b; Luo, Xu, and Liu 2023; Fuggitti and Chakraborti 2023), Answer Set Programming (ASP) (Yang, Ishay, and Lee 2023), and other variations of the paradigm (Pan et al. 2023; Wang et al. 2023; Chen et al. 2023a). Preliminary PDDL-LLM works are found in (Miglani and Yorke-Smith 2020; Feng, Zhuo, and Kambhampati 2018; Simon and Muise 2021; Chalvatzaki et al. 2023). Currently, researchers are further exploring the nuances of varying pipelines to balance the effectiveness and limitations of LLMs in building such neuro-symbolic frameworks. Huang et al. (2024b) survey a limited amount of works to compose their high-level abstraction of LLM-augmented planning agents. Pallagani et al. (2024) go beyond the scope of traditional AP, encapsulating more broader constructs, whereas Zhao et al. (2024) provide an extensive overview of LLM-TAMP applications. Most akin to our paper, Li et al. (2024a) and Wei et al. (2025) review studies using LLMs in planning with PDDL; however, their survey mainly consists of frameworks comprised of LLMs-as-Planners, rather than LLM-as-Modelers (cf. Figure 1).
@@
Scope of Survey: Broadly, LLMs+AP paradigms can be categorized as: (i) LLMs-as-Heuristics, where they enhance search efficiency by providing heuristic guidance (Silver et al. 2022; Tuisov, Vernik, and Shleyfman 2025; Sel, Jia, and Jin 2025); (ii) LLMs-as-Planners, either directly involved with the decision of action sequences (Zhang et al. 2024c) or proposing plans that are later refined through post-hoc implementations (Gundawar et al. 2024; Arora and Kambhampati 2023; Pallagani et al. 2022; Burns, Hughes, and Sycara 2024); and (iii) LLMs-as-Modelers, where LLMs assist in constructing planners (Katz et al. 2024; Cao et al. 2024) or planning models—defining planning domains, extracting task representations, and refining problem formulations. Our paper surveys approximately 80 papers and technical reports of the latter, focusing on key research questions and potential directions for future work in the context of LLMs and planning models.
Figure 2: Taxonomy of research in LLM Model Construction
3 Large Language Models for Constructing Planning Models
@@
We consider the research on leveraging powerful LLMs to assist in constructing planning models to be of critical importance and a more promising approach than relying on LLMs to perform planning directly (Vyas et al. 2025). Verifiable planning modules remain the backbone of planning, ensuring reliability, robustness, and explainability—core requirements for many planning tasks. Note that deploying LLMs themselves in an end-to-end manner to perform planning still falls short of providing soundness guarantees (Valmeekam et al. 2024) and inherently struggle with correctness and completeness (Katz et al. 2024). We organize the existing works into a taxonomy comprising three key areas: Model Generation, Model Editing, and Model Benchmarks, as shown in Figure 2. These tasks need joint efforts from the NLP and automated planning community.
@@ 3.1 Model Generation
This first group of works are centered around Model Generation—extracting components or generating complete planning models from the user or environmental natural language input. This is further divided into three aspects: Task Modeling (sec. 3.1.1), which translates objectives into initial conditions and goal states; Domain Modeling (sec. 3.1.2), which defines the foundational components like entities, actions, and relationships in the system; and Hybrid Modeling (sec. 3.1.3), which integrates both aspects to create a complete model, enabling end-to-end planning.
To facilitate further discussion on LLM-driven model construction, we highlight two key research questions:
  1. RQ1
    How can LLMs accurately align with human goals, ensuring these models correctly represent desired specifications and objectives?
  2. RQ2
    To what extent and granularity of detail can human-level instructions be effectively translated into accurate model definitions?
@@ 3.1.1 Task Modeling
For goal-only specification, Collins et al. (2022) and Grover and Mohan (2024) utilize few-shot prompting whereas Faithful CoT (Lyu et al. 2023) puts heavy emphasis on an interleaving technique of chain-of-thought (CoT) prompting (Wei et al. 2022b). Xie et al. (2023) assess the effectiveness of LLMs in translating tasks with varying levels of ambiguity in both NL and other languages such as Python. Text2Motion (Lin et al. 2023) leverage LLMs to generate goal predictions that are checked against a learned library of Q-function skills—based on feasibility—that support geometric feasibility in robotic tasks. Kwon, Kim, and Kim (2024) decomposes long-term tasks into sub-goals using LLMs, then executing task planning for each sub-goal with either symbolic methods or MCTS-based LLM planners. Safe Planner (Li et al. 2024b) uses an LLM to convert NL instructions into PDDL goals, enabling a closed-loop VLM-planner to operate based on real-time environmental observations. Fine-Morris et al. (2024) propose a framework that decomposes NL goals into predicate-based specifications, structured as Python dictionaries, which are then formatted into HDDL decomposition methods.
@@
Branching to multi-agent goal collaboration, DaTAPlan (Singh et al. 2024) employs an LLM to predict high-level anticipated tasks against human actions, triggering re-planning or new task predictions when deviations occur. PlanCollabNL (Izquierdo-Badiola et al. 2024) allocates sub-goals among LLM agents which are then encoded into PDDL goals, and modifying the action costs based on LLM recommendations. TwoStep (Singh, Traum, and Thomason 2024) decomposes multi-agent planning problems into two single-agent problems with mechanisms to ensure smooth coordination. LaMMA-P (Zhang et al. 2024d) allocate sub-tasks composed of general action sequences, and generates PDDL problem descriptions for each robot’s domain.
@@
Frameworks that handle complete PDDL task specifications can be broadly categorized into open-loop and closed-loop approaches. For the former, LLM+P (Liu et al. 2023a) uses in-context examples of problems within the same domain to generate whole problem files. TIC (Agarwal and Sreepathy 2024) achieved nearly 100% accuracy with GPT-3.5 Turbo across LLM+P planning domains by translating the task into intermediate representations, refining them, and processing them through a logical reasoner. (Lee et al. 2024) address a critical real-world application by introducing PlanAID, a system designed to generate emergency operation plans (EOPs) by leveraging Retrieval-Augmented Generation (RAG) documents to enhance the interaction between the user and the LLM. For real-world grounding, recent work converts NL instructions into structured geometric representations that bridge abstract language understanding and spatial reasoning (Ding et al. 2023b; Chen et al. 2023a; Chang et al. 2023), while other approaches integrate LLMs with Vision Language Models (VLMs) to further ground language understanding in spatial contexts (Shirai et al. 2023; Wang et al. 2025).
@@
For closed-loop approaches, Auto-GPT+P (Birr et al. 2024) generates the initial state of the problem based on visual perception and an automated error self-correction loop for the generated PDDL goal. Guo, Kingston, and Kavraki (2024) decompose the problem into both PDDL and Python specifications, incorporating a set of constraints into an SMT-based TAMP solver via a Python API. LLM+DP (Dagan, Keller, and Lascarides 2023) holds beliefs in uncertain environments to construct possible world states, dynamically updating its internal state and re-plans. PDDLEGO (Zhang et al. 2024a; Zhang 2024) performs a recursive task decomposition into sub-goals that enable the agent to gather new observations, progressively refining the problem file until it can develop a solvable plan. Liu et al. (2024c) integrate user information into a hierarchical scene graph of the environment, enabling an LLM to predict human activities and goal states, which are then refined using predicates and domain knowledge to ground problem specifications. Merino and Sabater-Mir (2024) model NPC behavior by leveraging LLMs conditioned on “memories” that represent environmental context.
@@
Instead of extracting PDDL models directly, Paulius et al. (2024) introduce a modularized approach that leverages LLMs to produce Object-Level Plans (OLP), which describe high-level changes to object states and uses them to bootstrap TAMP hierarchically. TRIP-PAL (de la Rosa et al. 2024) translates intermediate representations via travel points of interest (POI) and user information into dictionaries, which are then converted into PDDL planning tasks.
@@
Summary: Some methods directly translate natural language task descriptions into PDDL (Kelly et al. 2023; Liu et al. 2023a). Others enhance goal specification by incorporating reasoning chains and few-shot examples (Lyu et al. 2023), while (Dagan, Keller, and Lascarides 2023; Zhang et al. 2024a) target dynamic, online-planning applications. However, these current approaches rely on an explicit mapping between NL and PDDL code, requiring users to specify each state predicate in detail. To address ambiguity in minimal task descriptions, future research should develop methods capable of inferring complete and robust PDDL specifications from sparse input, building on prior work that has explored this concept via external perceptual groundings (Shirai et al. 2023), RAG implementations (Lee et al. 2024), or leveraging LLM commonsense capabilities to capture underlying assumptions and constraints (Agarwal and Sreepathy 2024).

### v2

# LLMs as Planning Formalizers: A Survey for Leveraging Large Language Models to Construct Automated Planning Models

@@ Abstract
Large Language Models (LLMs) excel in various natural language tasks but often struggle with long-horizon planning problems requiring structured reasoning. This limitation has drawn interest in integrating neuro-symbolic approaches within the Automated Planning (AP) and Natural Language Processing (NLP) communities. However, identifying optimal AP deployment frameworks can be daunting and introduces new challenges. This paper aims to provide a timely survey of the current research with an in-depth analysis, positioning LLMs as tools for formalizing and refining planning specifications to support reliable off-the-shelf AP planners. By systematically reviewing the current state of research, we highlight methodologies, and identify critical challenges and future directions, hoping to contribute to the joint research on NLP and Automated Planning.
LLMs as Planning Formalizers: A Survey for Leveraging Large Language Models to Construct Automated Planning Models
Marcus Tantakoun1,2, Christian Muise1,2, Xiaodan Zhu1,3 1Ingenuity Labs Research Institute, Queen’s University 2School of Computing, Queen’s University 3Department of Electrical and Computer Engineering, Queen’s University {20mt1, christian.muise, xiaodan.zhu}@queensu.ca
@@ 1 Introduction
The advent of Large Language Models (LLMs) has marked a significant paradigm shift in AI, sparking claims regarding emergent reasoning capabilities within LLMs (Wei et al., 2022a) and their potential integration into automated planning for agents (Pallagani et al., 2023). While LLMs, due to the prowess of distributed representation and learning, excel at System I tasks, planning—an essential aspect of System II cognition (Daniel, 2017)—remains a significant bottleneck (Bengio, 2020). Furthermore, LLMs face challenges with long-term planning and reasoning, often producing unreliable plans (Valmeekam et al., 2024b; Pallagani et al., 2023; Momennejad et al., 2023), frequently failing to account for the effects and requirements of actions as they scale (Stechly et al., 2024), with performance degrading under self-iterative LLM feedback (Stechly et al., 2023; Valmeekam et al., 2023b; Huang et al., 2024a).
@@
Figure 1: Distinction of planning using LLMs: (a) LLM-as-Planner uses LLMs for direct I/O planning; (b) LLM-as-Formalizer generates planning specifications for existing task planning methods (i.e. PDDL).
State-of-the-art LLMs have shown limited planning capability by directly generating action sequences—as the correctness, optimality, and reliability of their outputs are not guaranteed. Classical Automated Planning (AP) synthesizes plans through structured representation, logic, and search methods which are not subject to the weaknesses mentioned. Meanwhile, LLMs possess promising capabilities at extracting, interpreting, and refining planning model specifications from natural language (NL), acting as complementary components that can enable classical planners to generate robust solutions. Bringing the advantages of AP and LLMs together—LLMs for constructing planning specifications, AP systems for execution—defines the focus of this survey, a paradigm we call LLMs-as-Formalizers for constructing AP models.
@@
This paper is driven by the fragmented landscape in the current literature, where many surveys lack a cohesive overview of LLM integration within this field. Our focus stems from the need to address these gaps and provide a clear framework, highlighting the importance of aligning LLM capabilities with areas where they offer tangible benefits. The motivation for this approach is threefold: (i) Planning Accuracy: LLMs can help ensure that all relevant factors are considered, reducing the risk of overlooked constraints (Huang et al., 2024b). (ii) Adaptability: LLMs can aid systems in adapting to dynamic environments to capture real-world nuances better, reducing the need for manual edits (Dagan et al., 2023). (iii) Agnostic Modeling: LLMs trained on large corpora of diverse data can generalize across various domains without needing task-specific tuning, which reduces the reliance on specialized expertise, a major bottleneck in real-world AP integration (Gestrin et al., 2024).
@@
Our survey’s taxonomy is divided into three key areas: Model Generation, the process of extracting natural language input—originating from users or the environment—into structured planning model formalizations, further consisting of (i) Task Generation (sec. 3.1.1), which translates initial and goal states along with object assignments; (ii) Domain Generation (sec. 3.1.2), constructing predicates and action schemas; and (iii) Hybrid Generation (sec. 3.1.3), encapsulating both task instance and domain generation. Model Editing (sec. 3.2) systematically addresses code refinement and the repair of errors and inconsistencies in ill-defined planning formalizations. Finally, Model Benchmarks (sec. 3.3) encompass assessments of both the performance of LLMs in planning tasks and the quality of LLM-generated planning formalizations.
@@
To our knowledge, this is the first comprehensive survey of LLM-driven AP-model specification. Our contributions can be summarized as follows:
@@
  • A critical survey of LLM-driven Automated Planning (AP) model generation, editing, and AP-LLM benchmarks, structured within our taxonomy for a comprehensive field overview.
@@ 2.3 Large Language Models + Planning
The advancements of Large Language Models (LLMs) have shown promise in generating highly structured outputs, such as executable code, from NL descriptions (Li et al., 2023; Wang et al., 2024; Nijkamp et al., 2023). PDDL-LLM research is recent, with initial studies in (Miglani and Yorke-Smith, 2020; Feng et al., 2018; Simon and Muise, 2021; Chalvatzaki et al., 2023). Currently, researchers are further exploring the nuances of varying pipelines to balance the effectiveness and limitations of LLMs in building such neuro-symbolic frameworks. Huang et al. (2024c) survey a limited amount of papers to compose their high-level abstraction of LLM-augmented planning agents. Pallagani et al. (2024) go beyond the scope of traditional AP, encapsulating broader constructs, whereas Zhao et al. (2024) provide an extensive overview of LLM-TAMP applications. The work most akin to our paper is (Li et al., 2024a), which reviews studies using LLMs in planning with PDDL; however, their survey mainly consists of works comprising LLMs-as-Planners (cf. Figure 1).
@@
Scope of Survey: LLMs+AP is an expansive field encompassing many research areas, making it impractical to cover its entirety in a single survey with details and insights. Broadly, LLMs+AP paradigms can be categorized as: (i) LLMs-as-Heuristics, where LLMs enhance search efficiency via heuristic guidance (Silver et al., 2022; Hirsch et al., 2024; Tuisov et al., 2025; Sel et al., 2025); (ii) LLMs-as-Planners, where they either directly analyze action sequences (Zhang et al., 2024c; Lin et al., 2023) or propose plans that are refined through post-hoc methods (Gundawar et al., 2024; Arora and Kambhampati, 2023; Pallagani et al., 2022; Burns et al., 2024). In contrast, our survey analyzes (iii) LLMs-as-Formalizers, a paradigm in which LLMs are leveraged to construct AP models. Specifically, LLMs assist in defining planning model specifications, supported by domain-independent planners to generate solutions. Our paper surveys approximately 80 existing works, which utilize LLMs to construct planning models, discussing research questions that drive potential directions.
[…]
LLM Model Construction3) Model Benchmarks 3.3) Kambhampati et al. (2024);  Shridhar et al. (2021);  Guan et al. (2023);  Xie et al. (2024a);  Zheng et al. (2024);  Valmeekam et al. (2023a);  Kokel et al. (2024);  Stein and Koller (2023);  Bohnet et al. (2024);  Puerta-Merino et al. (2025);  Zuo et al. (2024);  Hu et al. (2025) Model Editing3.2) Gragera and Pozanco (2023);  Patil (2024);  Sikes et al. (2024a);  Caglar et al. (2024) Model Generation 3.1) Hybrid Modeling (§3.1.3) Kelly et al. (2023);  Smirnov et al. (2024);  Zhou et al. (2023);  Sakib and Sun (2024);  Liu et al. (2024b);  Huang and Zhang (2024);  Gestrin et al. (2024);  da Silva et al. (2024);  Hao et al. (2024);  Ying et al. (2023);  Ye et al. (2024);  Han et al. (2024);  Mahdavi et al. (2024);  Athalye et al. (2024);  Hu et al. (2024) Domain Modeling (§3.1.2) Oates et al. (2024);  Zhang et al. (2024b);  Sinha (2024);  Huang et al. (2024b);  Guan et al. (2023);  Ishay and Lee (2025);  Shah (2024);  Liu et al. (2024a);  Wong et al. (2023);  Ding et al. (2023a);  Chen et al. (2024);  Xie et al. (2024b);  Sikes et al. (2024b);  Oswald et al. (2024) Task Modeling (§3.1.1) Collins et al. (2022);  Grover and Mohan (2024);  Lyu et al. (2023);  Xie et al. (2023);  Kwon et al. (2024);  Li et al. (2024b);  Singh et al. (2024b);  Izquierdo-Badiola et al. (2024);  Singh et al. (2024a);  Zhang et al. (2024d);  Liu et al. (2023a);  Agarwal and Sreepathy (2024);  Agarwal et al. (2024);  Kalland (2024);  Ding et al. (2023b);  Chen et al. (2023a);  Chang et al. (2023);  Shirai et al. (2023);  Wang et al. (2025);  Birr et al. (2024);  Guo et al. (2024);  Dagan et al. (2023);  Zhang et al. (2024a);  Zhang (2024);  Lee et al. (2024);  Liu et al. (2024c);  Merino and Sabater-Mir (2024);  Paulius et al. (2024);  de la Rosa et al. (2024);  Fine-Morris et al. (2024)
Figure 2: Taxonomy of research in LLM Planning Model Specification
3 LLMs for Constructing Automated Planning Models
@@
We consider the research on leveraging powerful LLMs to assist in constructing planning models to be of critical importance. Verifiable planning modules remain the backbone of planning, ensuring reliability, robustness, and explainability. Note that deploying LLMs themselves in an end-to-end manner to perform planning still falls short of providing soundness guarantees (Valmeekam et al., 2024a) and may have principled weaknesses. We organize the existing works into a taxonomy comprising three key areas: Model Generation, Model Editing, and Model Benchmarks—where the term model, in this context, refers to AP specifications such as PDDL, as illustrated in Figure 2. The tasks require joint efforts from the NLP and automated planning community.
@@ 3.1 Model Generation
A large portion of this survey focuses on Model Generation—extracting and formalizing planning specifications from the user or environment via natural language input. This is further divided into three aspects: Task Modeling (sec. 3.1.1) defines objectives as initial conditions and goal states; Domain Modeling (sec. 3.1.2) defines the foundational components like entities, actions, and relationships in the system; and Hybrid Modeling (sec. 3.1.3) integrates both aspects to create a complete model, enabling end-to-end planning. A summary of core frameworks is provided in Appendix A.
To facilitate further discussion on LLM-driven planning model specification, we highlight two key research questions:
  1. RQ1:
    How can LLMs accurately align with human goals, ensuring these planning model specifications correctly represent desired expectations and objectives?
  2. RQ2:
    To what extent and granularity of detail can NL instructions be effectively translated into accurate planning model definitions?
@@ 3.1.1 Task Modeling
For goal-only specification, Collins et al. (2022) and Grover and Mohan (2024) utilize few-shot prompting whereas Faithful CoT (Lyu et al., 2023) puts heavy emphasis on an interleaving technique of chain-of-thought (CoT) prompting (Wei et al., 2022b). Xie et al. (2023) assess the effectiveness of LLMs in translating tasks with varying levels of ambiguity in both NL and other languages such as Python. Kwon et al. (2024) decompose long-term tasks into sub-goals using LLMs, then executing task planning for each sub-goal with either symbolic methods or MCTS-based LLM planners. Safe Planner (Li et al., 2024b) uses an LLM to convert NL instructions into PDDL goals, enabling a closed-loop VLM-planner to operate based on real-time environmental observations.
@@
Branching to multi-agent goal collaboration, DaTAPlan (Singh et al., 2024b) employs an LLM to predict high-level anticipated tasks against human actions, triggering re-planning or new task predictions when deviations occur. PlanCollabNL (Izquierdo-Badiola et al., 2024) allocates sub-goals among agents which are then encoded into PDDL, translating LLM output sub-goals into PDDL goals, and modifying the action costs based on LLM recommendations. TwoStep (Singh et al., 2024a) decompose multi-agent planning problems into two single-agent problems with mechanisms to ensure smooth coordination. LaMMA-P (Zhang et al., 2024d) uses LLMs to allocate sub-tasks composed of general action sequences, and generates PDDL problem descriptions for each robot’s domain.
@@
Frameworks that handle complete PDDL task specifications can be broadly categorized into open-loop and closed-loop approaches. In open-loop systems, LLM+P (Liu et al., 2023a) uses in-context (IC) examples of a related NL problem and its PDDL representation to generate whole problem files. TIC (Agarwal and Sreepathy, 2024) achieved nearly 100% accuracy with GPT-3.5 Turbo across LLM+P planning domains by translating the task into intermediate representations, refining them, and processing them through a logical reasoner. Kalland (2024) combines a language and an automatic speech recognition model to generate PDDL instances. Recent work converts NL instructions into structured geometric representations that bridge abstract language understanding and spatial reasoning for task and motion planning (Ding et al., 2023b; Chen et al., 2023a; Chang et al., 2023), while other approaches integrate LLMs with Vision Language Models (VLMs) for visual perception to further ground language understanding in spatial contexts (Shirai et al., 2023; Wang et al., 2025).
@@
In closed-loop systems, Auto-GPT+P (Birr et al., 2024) generates the initial state of the problem based on visual perception and an automated error self-correction loop for the generated PDDL goal. Guo et al. (2024) decompose the problem into both PDDL and Python specifications, incorporating a set of constraints into an SMT-based TAMP solver via a Python API. LLM+DP (Dagan et al., 2023) holds beliefs in uncertain environments to construct possible world states, dynamically updating its internal state and re-plans. PDDLEGO (Zhang et al., 2024a; Zhang, 2024) performs a recursive task decomposition into sub-goals that enable the agent to gather new observations, progressively refining the problem file until it can develop a solvable plan. In terms of human-in-the-loop collaboration, (Lee et al., 2024) introduce PlanAID, a system that uses Retrieval-Augmented Generation (RAG) to assist LLMs in generating emergency operation plans (EOPs) through improved user interaction. Liu et al. (2024c) integrate user information into a hierarchical scene graph of the environment, enabling an LLM to predict human activities and goal states, which are then refined using predicates and domain knowledge to ground problem specifications. Merino and Sabater-Mir (2024) model NPC behavior by leveraging LLMs conditioned on “memories” that represent environmental context.
@@
Rather than extracting PDDL directly, Paulius et al. (2024) leverages LLMs to produce Object-Level Plans (OLP), which describe high-level changes to object states and uses them to bootstrap TAMP hierarchically. TRIP-PAL (de la Rosa et al., 2024) translates intermediate representations via travel points of interest (POI) and user information into dictionaries. Fine-Morris et al. (2024) decompose NL goals into predicate-based Python dictionaries, which are then formatted into HDDL decomposition methods. Beyond PDDL, LLM+AP has expanded generating other planning specifications (Pan et al., 2023) such as temporal logic (TL) representations (Chen et al., 2023b), including TSL (Murphy et al., 2024), STL (Mao et al., 2024), and LTL (Cosler et al., 2023; Liu et al., 2023b; Manas et al., 2024; Luo et al., 2023). Recent work shows that LLMs can define the task search space by generating successor and goal state Python functions, enabling classical solvers to explore the space efficiently (Katz et al., 2024; Cao et al., 2024).
@@
Summary and Future Directions: Some methods directly translate NL task descriptions into PDDL (Kelly et al., 2023). Others enhance goal specification by incorporating reasoning chains and few-shot examples (Lyu et al., 2023; Liu et al., 2023a). However, these current approaches rely on an explicit mapping between NL and PDDL code, limiting their processes as code translation tasks. To address ambiguity in minimal task descriptions, future research should develop methods capable of inferring complete and robust PDDL specifications from sparse input, building on prior work that has explored this concept via external perceptual groundings (Shirai et al., 2023), RAG implementations (Lee et al., 2024), or leveraging LLM commonsense capabilities to capture underlying assumptions and constraints (Agarwal and Sreepathy, 2024).